Can you help me simplify the expression first by using the distributive property to remove any parentheses.
-7(z - 4)
Can you solve these equations
1.) 5z -(3x - 1)= 2
2.) 1/3x - 1/3 = -2
3.) 1/3x - 4 = 1

Answer by stanbon(75887)

Can you help me simplify the expression first by using the distributive property to remove any parentheses.
-7(z - 4) = -7z+28
----------------------------
Can you solve these equations
1.) 5z -(3x - 1)= 2
5z-3x+1 = 2
5z-3x = 1
Comment: Can't solve because you have two variables.
--------------------------------------------------------
2.) 1/3x - 1/3 = -2
(1/3)x = -2 + (1/3)
(1/3)x = -(6/3)+(1/3)
(1/3)x = -5/3
x = -5
--------------------------
3.) 1/3x - 4 = 1
Multiply thru by 3 to get:
x -12 = 3
x = 15
==================
Cheers,
Stan H.